f(x).f(y)=f(x+y)
find f(x)

Grade:10

2 Answers

Arun Kumar IIT Delhi
8 years ago
Hello Student,


\\f(x)f(y)=f(x+y) \\=>f'(x)f(y)=f'(x+y)---eq1 \\y=0
\\=>f(0)=1 \\=>f(x)f(-x)=1----eq2 \\$from 1 & 2$ \\f(x)=e^x
Arun Kumar IIT Delhi
8 years ago
\\f(x)f(y)=f(x+y) \\=>f'(x)f(y)=f'(x+y)---eq1 \\y=0 \\=>f(0)=1 \\=>f(x)f(-x)=1----eq2 \\$from 1 & 2$ \\f(x)=e^x